Question 5 Tries remaining: 2 Points out of 7.70 The demand for okra is given by: Q 220 -P. The supply of okra is given by: Qs 8P-50. The government has implemented a price floor of $58. Calculate producer surplus with the price floor (Do not include a $ sign in your response. Round to the nearest two decimal places if necessary.) Answer: Check
Answer #1

Demand is given by Q = 220 - P

Supply is given by Q = 8P - 50

Market equilibrium has quantity demanded and supplied equal

220 - P = 8P - 50

270 = 9P

P = 30 and Q = 190.

At a price floor of 58, market quantity us 220 - 58 = 162. The price that sellers are willing to receive at this quantity is 162 = 8P - 50 or P = 26.50

Producer surplus = (price floor - price that sellers are willing to receive at market quantity)*(market quantity) + 0.5*(price that sellers are willing to receive at market quantity - minimum acceptable price)*market quantity

= (58 - 26.50)*(162) + 0.5*(26.50 - 6.25)*162

= $6743.25